OpenStudy (anonymous):

For the Integral (1/X^p) dx from 0 to 1, find the values of p for which the integral converges ?

OpenStudy (anonymous):

i can integrate it don't know about convergence

OpenStudy (anonymous):

\[ \int _0^1 x^ {-p} dx\] \[ \left| \frac{x^{1-p}}{1-p}\right|_0^1\]

OpenStudy (anonymous):

\[1 - p \neq 0\] \[p \neq 1 \]